### v1

# Particle motion around luminous neutron stars: effects of deviation from Schwarzschild spacetime

Abstract
We study trajectories of test particles around a luminous, static, spherically symmetric neutron star, under the combined influence of gravity and radiation. In general relativity, for Schwarzschild spacetime, an equilibrium sphere (the Eddington Capture Sphere) is formed for near-Eddington luminosities. We generalize these results to a broad class of static, spherical spacetimes. We also study the dynamics of particles in a strong radiation field in spherical spacetimes. The results are illustrated for two cases, Reissner-Nordström spacetime of a charged spherical object in general relativity and Kehagias-Sfetsos spacetime, arising from the Hořava-Lifshitz gravity theory. Our findings apply to neutron stars under gravitational field equations different from the vacuum Einstein field equations of general relativity, such as in modified theories of gravity, the only requirement being that test particles follow geodesics in the absence of the radiation field. We show that it is possible, in principle, to test the gravity theory by investigating such test particle equilibria and trajectories with observations of neutron stars’ X-ray bursts.
I Introduction
Radiation exerts force on test particles, influencing their trajectories and causing their deviation from timelike geodesics. For a luminous source the radiation force may be comparable to the effective gravity. In particular, the Eddington luminosity
$L_{\rm{Edd}}=\frac{4\pi mGMc}{\sigma_{\rm T}}$ (1)
results in a global equilibrium between the Newtonian gravity and the radiation pressure for a spherical source of mass $M$. In Eq. (1) we assume hydrogen plasma, hence $\sigma_{\rm T}$ is the Thomson cross section and $m$ is the proton mass. $G$ and $c$ denote the gravitational constant and speed of light, respectively. When it is not confusing, we use $G=c=1$ units hereafter. The problem of the gravity-radiation equilibrium is more complicated for the relativistic case – in general relativity (GR) equilibrium can only be established at a particular radius, referred to as the Eddington Capture Sphere (ECS; see Stahl2012 ; Wielgus2012 ). The ECS radius $r_{\rm ECS}$ depends on the stellar luminosity as seen by a distant observer $L_{\infty}$, and in Schwarzschild spacetime corresponds to
$r_{\rm ECS}=\frac{2\,GM/c^{2}}{1-\lambda^{2}}\ ,$ (2)
where $\lambda\equiv L_{\infty}/L_{\rm Edd}$. In Eq. (2), $L_{\infty}=L_{\rm{Edd}}$ represents equilibrium established at infinity. The ECS may form as a transient phenomenon during type-I luminous X-ray bursts on neutron stars Tawara1984 ; Lewin1993 , particularly in cases exhibiting photospheric expansion Lewin1984 ; Strohmayer2006 . Studying such events may lead to constraints on neutron star equation of state Ozel2006 ; Lattimer2012 , mass, radius Galloway2008 ; Bollimpalli2019 , and spin Muno2001 . In this paper we consider whether the location of the equilibrium surface and test-particle dynamics are sensitive to deviations of the underlying metric from the Schwarzschild solution. While several authors considered deviations caused by the central mass rotation SokOh2010 ; Bini2011 ; DeFalco2019 ; Wielgus2019 , quadrupolar deviation of the metric from spherical symmetry DeFalco2020 , or both DeFalco2021 , we focus on the effects related to modifying the underlying gravity field equations (and/or modifications of the background metric).
Furthermore, the aim of this paper is to demonstrate the impact of the spacetime deviation from the canonical Schwarzschild case on properties of systems including luminous neutron stars, illustrating it with concrete examples. Such deviations in spacetime geometry could be in principle detected through detailed modeling of X-ray light curves of bursting neutron stars, offering a previously unexplored avenue to test theories of gravity in the strong field and large curvature regime.
II Equations of motion
We consider a static, spherically symmetric spacetime metric in the form
${\rm d}s^{2}=-\xi\,c^{2}{\rm d}t^{2}+\xi^{-1}\,{\rm d}r^{2}+r^{2}\,{\rm d}\theta^{2}+r^{2}\sin^{2}\theta\,{\rm d}\phi^{2}\ ,$ (3)
which describes a wide class of spherically symmetric spacetimes, where $\xi$ depends only on the radial coordinate $r$, $\xi\equiv\xi(r)$. Schwarzschild spacetime is recovered for $\xi(r)=1-2M/r$, where $M$ is the geometric mass of the star. In the presence of a radiation flux $F^{\mu}$, the test-particle equation of motion reads
$a^{\mu}=u^{\nu}\nabla_{\nu}u^{\mu}=\frac{\rm d^{2}x^{\mu}}{\rm d\,{\tau}^{2}\,}+\Gamma^{\mu}_{\ \nu\rho}u^{\nu}u^{\rho}=\frac{\sigma_{\rm T}}{m{c}}F^{\mu}\ ,$ (4)
where $a^{\mu}$ is the particle’s four-acceleration and $u^{\mu}$ is its four-velocity. The only non-zero Christoffel symbols are
$\Gamma^{r}_{\ tt}=\frac{c^{2}\xi\xi^{\prime}}{2}\ ,$ (5)
$\Gamma^{r}_{\ rr}=-\frac{\xi^{\prime}}{2\xi}\ ,$ (6)
$\Gamma^{r}_{\ \theta\theta}=-r\xi\ ,$ (7)
$\Gamma^{r}_{\ \phi\phi}=-r\xi\sin^{2}\theta\ ,$ (8)
$\Gamma^{\phi}_{\ r\phi}=\frac{1}{r}\ ,$ (9)
where we denote
$\xi^{\prime}\equiv\frac{{\rm d}\xi}{{\rm d}r}\ .$ (10)
Using the projection tensor $h^{\mu}_{\ \nu}$ onto the particle’s local 3-space, Kronecker tensor $\delta^{\mu}_{\ \nu}$, and the radiation stress-energy tensor $T^{\nu\rho}$, the radiation flux $F^{\mu}$ measured in the particle’s local rest frame can be represented as
$F^{\mu}=h^{\mu}_{\ \nu}T^{\nu\rho}u_{\rho}=\left(\delta^{\mu}_{\ \nu}+u^{\mu}u_{\nu}\right)T^{\nu\rho}u_{\rho}\ .$ (11)
A fully covariant description of the radiation field of a static, uniformly radiating, spherical star in Schwarzschild spacetime in terms of the radiation tensor $T^{\nu\rho}$ was first given by Abramowicz1990 . This description can be readily generalized to a broader class of static spherical metrics. The frequency-integrated specific intensity $I$ of radiation is uniform over all angles and given by
$I(r)=I(R)\left(\frac{\xi_{R}}{\xi}\right)^{2}\,,$ (12)
where $R$ is the stellar radius and $\xi_{R}\equiv\xi(R)$. The apparent viewing angle $\alpha_{0}$ of the star, measured at radius $r>R$, is
$\sin\alpha_{0}=\frac{R}{r}\left(\frac{\xi}{\xi_{\rm R}}\right)^{1/2}\ .$ (13)
Equation (13) naturally requires that $r>R$, with another requirement that $r>r_{\gamma}$, where $r_{\gamma}$ is the photon sphere radius, a solution to $\xi^{\prime}r-2\xi=0$, see e.g. Wielgus2021 . It is generally expected that $R>r_{\gamma}$ for neutron stars, while it may not necessarily be the case for hypothetical more compact strange stars, e.g. Haensel1986 .
Figure 1: Radius of gravity-radiation equilibrium sphere for Reissner-Nordström (top row) and Kehagias-Sfetsos (bottom row) spacetimes for various luminosity parameters $\lambda$. Left column: radius as function of the metric parameter $(Q/M)^{2}$ in RN spacetime ($\omega M^{2}$ in KS spacetime), $Q=0$ ($\omega M^{2}\rightarrow\infty$) corresponds to Schwarzschild spacetime. Right column: relative change in the location of the equilibrium sphere with respect to the Schwarzschild case.
Figure 2: Radius (left column), radial velocity component $v^{r}=\sqrt{-g_{rr}/g_{tt}}\,u^{r}/u^{t}$ (middle column), and energy $u_{t}$ (right column) as functions of the proper time $\tau$ for particles ejected from the neutron star surface located at $R=4M$ by the instantaneous burst of radiation with the luminosity parameter $\lambda=0.75$ (solid), $\lambda=0.80$ (dashed), and $\lambda=0.85$ (dotted). The black lines correspond to Schwarzschild spacetime. The red lines in the top row correspond to Reissner-Nordström spacetime with $(Q/M)^{2}=1$. The red lines in the bottom row correspond to Kehagias-Sfetsos spacetime with $\omega M^{2}=0.5$.
The radiation stress-energy tensor is given by $T^{\mu\nu}=I(r)\,A^{\mu\nu}$, where the dimensionless components of the $A^{\mu\nu}$ tensor take the following form in the orthonormal tetrad associated with the coordinate basis vectors,
$A^{\hat{t}\hat{t}}$ $=2\pi\left(1-\cos\alpha_{0}\right)\ ,$ (14)
$A^{\hat{t}\hat{r}}$ $=\pi\sin^{2}\alpha_{0}\ ,$ (15)
$A^{\hat{r}\hat{r}}$ $=\frac{2\pi}{3}\left(1-\cos^{3}\alpha_{0}\right)\ ,$ (16)
$A^{\hat{\phi}\hat{\phi}}$ $=A^{\hat{\theta}\hat{\theta}}=\frac{\pi}{3}\left(\cos^{3}\alpha_{0}-3\cos\alpha_{0}+2\right).$ (17)
Tetrad frame components $A^{\hat{\mu}\hat{\nu}}$ are related to the coordinate frame components $A^{\mu\nu}$ through
$A^{\hat{\mu}\hat{\nu}}=A^{\alpha\beta}e_{\ \alpha}^{\hat{\mu}}e_{\ \beta}^{\hat{\nu}}\,,$ (18)
where tetrad components in the coordinate basis are given by
$e_{\ \mu}^{\hat{\mu}}=|g_{\mu\mu}|^{1/2}.$ (19)
We parameterize the radiation strength in terms of the dimensionless parameter $\lambda$
$\lambda=\frac{L_{\infty}}{L_{\rm Edd}}=\frac{\xi_{R}L(R)}{L_{\rm Edd}}$ $=\frac{4\pi^{2}\xi_{R}R^{2}I(R)}{L_{\rm Edd}}$
$=\frac{\pi\sigma_{\rm T}\xi_{R}R^{2}I(R)}{GMmc}$ (20)
where $M$ in a general spacetime corresponds to the ADM mass, so we assume that the spacetime approaches Schwarzschild solution in the weak field (large radius) limit. Putting together Eqs. (4)–(20) and leveraging spherical symmetry to choose a coordinate system with $\theta=\pi/2$, ${\rm d}\theta/{\rm d}\tau={\rm d}^{2}\theta/{\rm d}\tau^{2}=0$, we obtain the full equations of motion for a test particle
$\frac{{\rm d}^{2}r}{{\rm d}\tau^{2}}$ $=$ $-\frac{1}{2}\xi^{\prime}+r\,\left(\xi-\frac{r\xi^{\prime}}{2}\right)\bigg{(}\frac{{\rm d}\phi}{{\rm d}\tau}\bigg{)}^{2}$
$+\frac{\lambda M\,\xi_{R}}{\pi R^{2}\xi^{2}}\,\left[\xi\,A^{\hat{t}\hat{r}}u^{t}-\left(A^{\hat{r}\hat{r}}+\chi\right)\frac{{\rm d}r}{{\rm d}\tau}\right]\ ,$
$\frac{{\rm d}^{2}\phi}{{\rm d}\tau^{2}}$ $=$ $-\left[\frac{1}{r}\frac{{\rm d}r}{{\rm d}\tau}+\frac{\lambda M\,\xi_{R}}{\pi R^{2}\xi^{2}}\,\left(A^{\hat{\phi}\hat{\phi}}+\chi\right)\right]\frac{{\rm d}\phi}{{\rm d}\tau}\ ,$ (22)
where $\tau$ is the particle’s proper time and we denote
$\chi\equiv$ $\xi\,A^{\hat{t}\hat{t}}(u^{t})^{2}+\xi^{-1}A^{\hat{r}\hat{r}}\left(\frac{{\rm d}r}{{\rm d}\tau}\right)^{2}$ (23)
$+\,r^{2}\,A^{\hat{\phi}\hat{\phi}}\left(\frac{{\rm d}\phi}{{\rm d}\tau}\right)^{2}-2A^{\hat{t}\hat{r}}u^{t}\,\frac{{\rm d}r}{{\rm d}\tau}.$
Additionally, from the four-velocity normalization $u^{\alpha}u_{\alpha}=-1$ we have
[…]
There are four components in Eq. (II), that can be associated with gravitational acceleration $\propto\xi^{\prime}$, centrifugal acceleration $\propto({\rm d}\phi/{\rm d}\tau)^{2}$, radiation pressure $\propto\lambda\,u^{t}$, and radial radiation drag $\propto\lambda\,{\rm d}r/{\rm d}\tau$. In Eq. (22) we recognize the Coriolis acceleration $\propto({\rm d}r/{\rm d}\tau)({\rm d}\phi/{\rm d}\tau)$. The second component of Eq. (22) is the Poynting–Robertson drag force, removing particle’s angular momentum through interaction with the radiation field. In general relativity these effects were studied in, e.g., Bini2009 ; Stahl2013 ; Mishra2014 ; DeFalco2019 ; Fragile2020 . Equations (II)–(23) reduce to the ones presented in Stahl2012 for the case of Schwarzschild spacetime.
II.1 Radial equilibrium surface
For a static test particle we have $u^{\mu}=\delta^{\mu}_{\ t}u^{t}$ and $u^{t}=\xi^{-1/2}$, hence the only non-zero components in Eq. (II) correspond to gravitational acceleration and radiation pressure and Eq. (22) is trivially fulfilled. The whole system of equations of motion simplifies to
$0=-\frac{1}{2}\xi^{\prime}+\frac{\lambda M}{r^{2}\xi^{1/2}}$ (25)
for $r>r_{\gamma}$ and at the equilibrium radius we have
$2M\lambda=r^{2}\xi^{1/2}\xi^{\prime}\,.$ (26)
If Eq. (26) has a solution for some $r=r_{0}>R$, then it physically corresponds to the gravity-radiation equilibrium surface suspended above the luminous stellar surface. For Schwarzschild metric Eq. (26) reduces to $\lambda=\xi^{1/2}$, and hence the ECS formula given in Eq. (2) is recovered. We denote this general equilibrium radius with $r_{0}$, and the Schwarzschild value $r_{\rm ECS}\equiv r_{0,\rm{Schw}}$.
To discuss the stability of the equilibrium, let us consider a restoring force
$f(r)=\frac{1}{2}\xi^{\prime}-\frac{\lambda M}{r^{2}\xi^{1/2}}\,,$ (27)
with $f(r_{0})=0$. In order for the radial equilibrium at $r_{0}$ to be stable, we need
$f^{\prime}(r)=\frac{1}{2}\xi^{\prime\prime}+\frac{\lambda M}{2r^{2}\xi^{1/2}}\left(\frac{\xi^{\prime}}{\xi}+\frac{4}{r}\right)$ (28)
to be positive at $r_{0}$. Using the $f(r_{0})=0$ condition we find the stability condition at the equilibrium radius
$f^{\prime}(r_{0})=\frac{1}{2}\xi^{\prime\prime}+\frac{1}{4}\xi^{\prime}\left(\frac{\xi^{\prime}}{\xi}+\frac{4}{r_{0}}\right)>0\ .$ (29)
Stability of gravity-radiation equilibria is thus a spacetime property, and for Schwarzschild spacetime equilibria are radially stable for all $r_{0}>r_{\gamma}$ Abramowicz1990 ; Stahl2012 . For the cases discussed further in this paper we always find stability for $r_{0}>r_{\gamma}$.
III Metric examples
We exemplify the general calculations given in Section II by considering two particular spacetime models: Reissner-Nordström (RN) and Kehagias-Sfetsos (KS; kehagiasSfetsos2009PhLB, ). The former originates from the electrovacuum Einstein field equations and corresponds to the spacetime of a static electrically charged black hole (or exterior of a charged spherical star). Recently a lot of interest in RN spacetime came from the brane cosmology Randall1999 , where the RN model is a solution of the effective field equation and the “tidal charge” corresponds to gravitational effects from the fifth dimension Dadhich2000 . The KS spacetime arises as a static, spherical solution representing a black hole (or exterior of a star) in Horava’s gravity, a popular model of a quantum field theory of gravity Horava2009 . The geodesic structure of the KS spacetime has been studied extensively by katkaVieiraEtal2015GRG ; stuchlikSchee2014CQG ; vieiraMarekEtal2014PRD , for instance.
Figure 3: Trajectories of particles ejected from the surface of a neutron star in Reissner-Nordström (top, with $(Q/M)^{2}=1$) and Kehagias-Sfetsos (bottom, with $\omega M^{2}=0.5$) spacetimes. We assume a stellar radius of $R=4M$ (gray disc) and a luminosity parameter $\lambda=0.85$. The continuous red circles represent the corresponding equilibrium spheres, while the dashed red circles correspond to the Schwarzschild equilibrium (ECS). Each pair of colored trajectories corresponds to a particle ejected with a different initial velocity with respect to a local static observer at the stellar surface (starting at the marked dots): $v^{r}=0.15$, $v^{\phi}=0.2$ (blue); $v^{r}=0$, $v^{\phi}=0.4$ (green); $v^{r}=0.05$, $v^{\phi}=0.3$ (magenta). Continuous curves correspond to trajectories in RN and KS spacetimes, respectively; dashed curves represent trajectories in Schwarzschild spacetime computed for the same initial conditions.
III.1 Equilibrium spheres
In RN spacetime we have
$\xi_{\rm RN}(r)=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}},$ (30)
where $Q$ is the star’s charge, and the radial equilibrium Eq. (26) gives a 4th order polynomial equation
$0=$ $\,M^{2}(1-\lambda^{2})r^{4}-2M(Q^{2}+M^{2})r^{3}$
$+Q^{2}(Q^{2}+5M^{2})r^{2}-4Q^{4}Mr+Q^{6}.$ (31)
The polynomial has a single real positive root $r_{0}$ such that $r_{0}>M$, corresponding to the equlibrium sphere. The value of this equilibrium radius $r_{0}$ as a function of charge and luminosity parameter $\lambda$ is explored in Fig. 1 (top row). We see from the Figure that for all considered values of $\lambda$, we have $r_{0}/r_{0,\rm Schw}>1$. The equilibrium sphere radius generally grows with the charge $Q$ and can significantly deviate from the Schwarzschild solution for large values of $(Q/M)$ and large luminosity.
In KS spacetime we have
$\xi_{\rm KS}(r)=1+\omega r^{2}\left[1-\left(1+\frac{4M}{\omega r^{3}}\right)^{1/2}\right].$ (32)
The parameter $\omega$ is such that for $\omega\to\infty$ we recover Schwarzschild spacetime. Here we evaluate the location of the equilibrium sphere by solving Eq. (26) numerically, and the results are presented in Fig. 1 (bottom row). The equilibrium sphere radius is once again larger than for the Schwarzschild case, but it is actually the lower neutron star luminosity (equilibrium forming for smaller $r_{0}$) resulting in a larger fractional deviation.
Figure 4: Hoyle-Lyttleton accretion onto a luminous star. The particles are initiated at horizontal axis coordinate $x=-5000M$, where the gravitational and radiative impact of the star is negligible, all with initial velocity of $v=0.1\,c$ along the $x$ direction. The star is characterized by $R=4M$ and $\lambda=0.99$. Left: Schwarzschild spacetime. Trajectories with relatively low impact parameter are captured by the equilibrium sphere, whereas trajectories with high impact parameter are scattered (to infinity). The red curve corresponds to the lowest impact parameter for scattered trajectories (to infinity) considering a $1M$ resolution in the grid of impact parameters, $b_{\rm crit}=1697M$. Middle: RN spacetime with $(Q/M)^{2}=1$. The red curve curve corresponds to the lowest impact parameter for scattered trajectories $b_{\rm crit}=1615M$. Typical captured trajectories in the middle panel penetrate the equilibrium sphere and then bounce, crossing it outwards and finally settling into it. The same qualitative picture happens in the right panel (KS spacetime with $\omega M^{2}=0.5$), but with a much smaller bouncing amplitude and with $b_{\rm crit}=1697M$, resembling the behavior in the Schwarzschild spacetime.
III.2 Particle ejection
We study the scenario of a luminous outburst, expelling particles from the neutron star surface. Under the assumption of spherical symmetry, this becomes a one-dimensional problem for the particle’s radial motion. Similar problems were considered for the Schwarzschild case, e.g., by Wielgus2012 ; Stahl2013 ; Mishra2014 , including more general configurations. Here, we consider only the most physically relevant case, corresponding to particles ejected (with zero initial velocity) from the neutron star surface by a sudden burst of luminosity to a near-Eddington value. This is essentially a toy model for the photospheric radius expansion X-ray bursts of neutron stars Lewin1984 ; Lewin1993 ; Kuulkers2003 . Example results are shown in Fig. 2. The timescale for the particles to settle on the equilibrium sphere is no longer than 1000 $M$ for the chosen peak luminosity $\lambda$, which is of order of 0.01 s for a typical neutron star. The radial velocity of ejected particles may be significantly larger for non-Schwarzschild spacetimes, reaching around 0.1 $c$ for RN spacetime, before the particle settles on the equilibrium sphere. The kinetic energy of the particle is removed by the radiative drag force (Fig. 2, second column), but the particle energy $u_{t}$ is not conserved and grows as a result of the radiation pressure (Fig. 2, last column).
In Fig. 3 we present orbits of particles ejected from the stellar surface with a sudden blast of luminosity for RN and KS spacetimes, with nonzero initial velocity components with respect to the star, comparing them with results for Schwarzschild spacetime obtained for the same initial condition. The figure demonstrates the efficiency of the angular momentum removal through radiation Poynting-Robertson drag force. The larger variation, related to detailed balance of gravity, radiation pressure and radiative drag forces, can be observed once again, particularly in the RN trajectories.
III.3 Hoyle-Lyttleton accretion
We consider a scenario of a luminous star passing through a cloud of matter with a fixed velocity, a Hoyle-Lyttleton accretion problem Hoyle1939 . This simple setup allows us to systematically explore the problem of the dynamics of particles with non-zero angular momentum with respect to the neutron star and the impact of the spacetime geometry.
In Fig. 4 we present trajectories of particles in the Hoyle-Lyttleton scenario for Schwarzschild (left), extreme RN (middle), and extreme KS (right) spacetimes. The setup corresponds to a collection of particles approaching the neutron star from the left with an initial velocity of 0.1 $c$, set beyond the gravitational sphere of influence of the neutron star. The particles are parameterized by their impact parameter, proportional to their (initial) specific angular momentum with respect to the neutron star. Particles with impact parameter $b$ lower than certain critical value $b_{\rm crit}$ (specific angular momentum $\ell<\ell_{\rm crit}$) are captured by the equilibrium sphere, hence the name Eddington Capture Sphere. The value of the $b_{\rm crit}$ parameter (and therefore of $\ell_{\rm crit}$) depends on the initial velocity, and also (weakly) on the neutron star diameter, impacting dynamical radiation drag forces and hence the efficiency of the angular momentum removal. These effects were investigated by, among others, Bini2009 ; Stahl2012 for the Schwarzschild spacetime case and by SokOh2011 ; Wielgus2019 for a slowly rotating neutron star. Here we demonstrate that the role of the spacetime geometry may also be significant, which is particularly clear for the (extreme) RN case in Fig. 4. It is interesting to observe that (extreme) RN spacetime, with the equilibrium sphere larger than those of Schwarzschild and (extreme) KS cases, actually corresponds to a smaller cross section for capturing particles. This highlights the dynamic character of this problem, and the role of the relativistic radiation drag forces acting on a moving particle and depending on its four-velocity.
IV Discussion and summary
We demonstrated the importance of the spacetime geometry for equilibria and dynamics of particles around luminous compact stars. In the astrophysical context, understanding the location of the gravity-radiation equilibrium surface is important for the observational constraints on the neutron star equation of state from the photospheric radius expansion X-ray bursts in the touchdown phase Ozel2006 ; Galloway2008 ; Kim2021 , or from oscillations of the material suspended above the star surface by the strong radiation Bollimpalli2019 . On the one hand, relaxing the Schwarzschid spacetime model assumption results in another source of uncertainty, on the other hand, it may offer a way to test strong gravity. What is more, the dynamical properties of such systems could manifest observationally, e.g., increased velocity of the expanding atmosphere of a bursting neutron star would necessarily manifest as a shift in the observed X-ray spectrum. These problems have not received sufficient attention from the scientific community so far.
There are multiple research avenues that should be followed in order to generalize the presented results and make them more physically relevant. In particular, the test-particle approximation can be relaxed and more realistic cases of optically thin Wielgus2015b ; Bollimpalli2017 and optically thick Wielgus2016 levitating (suspended in the radiation field) atmospheres can be considered. The questions regarding the impact of the rotation of the star and the role of magnetic fields are also important as subjects of further studies that will necessarily render the presented toy model more complicated.
Acknowledgements
We thank Marek Abramowicz and Włodek Kluźniak for discussions. We also thank Alexandra Elbakyan for her contributions to the open science initiative. MW acknowledges support from the European Research Council advanced grant “M2FINDERS - Mapping Magnetic Fields with INterferometry Down to Event hoRizon Scales” (Grant No. 101018682).